The number of sheep is 8/3 rd the number of kangaroos. Together they total 88,000. How many kangaroos are there? I need to write a formula and solve the problem. Thank you for any help! ( Desperate mom that never took algebra)
Found 2 solutions by dkppathak, Alan3354:
Answer by dkppathak(439)

The number of sheep is 8/3 rd the number of kangaroos. Together they total 88,000. How many kangaroos are there?
let the number of kangaroos are X
sheeps will be 8x/3
total sheep + kangaroo = 88
X+8x/3 =88
11x/3=88 by cross multiplication
11x=88x3
x=88x3/11=24
answer kangaroo are 24
sheeps are 88-24=64

Answer by Alan3354(69443)

8/3 means 8 thirds.
Don't add the rd to it.